OpenStudy (lilai3):

Could someone please help me with these 2 questions? A card is chosen at random from a deck of 52 cards. It is then replaced and a second card is chosen. What is the probability of getting a jack and then an eight? A standard deck of playing cards contains 52 cards in 4 suits of 13 cards each/ 2 suits are read and 2 suits are black. Find each probability. Assume the first card is replaced before the second card is drawn. P(black, queen) P(black, diamond) P(jack, queen)

OpenStudy (mathmagician):

1. Jack and 8: P=(4/52)*(4/52)=0.0059

OpenStudy (experimentx):

first:4/52x4/52 P(black, queen) = 1/2x(4/52) P(black, diamond) = 1/2x(13/52) P(jack, queen) = (4/52)x(4/52)

OpenStudy (anonymous):

sampling with replacement, you can assume the probabilities are the same (the replacement implies nothing has changed) so when you pick a jack, idk what that is but its probability is the amount of jacks/the total amout of cards. you do the same for the probability of the eight. in the second part you replace again and make the assumption that your draws are independent. this makes it so you can just multiply the odds of event A with event B to get the probability of both.

OpenStudy (anonymous):

There are thirteen kinds of cards. The probability of getting a Jack from a full deck is one in thirteen (4/52). We replace the Jack, and draw again. The probability of getting an eight is again one of thirteen. The combined probability is the product, one of 169.

OpenStudy (lilai3):

Oh, I get it now, guys. Thank you guys sooo much!!! :]

OpenStudy (lilai3):

oh yeah...animal ain i haf one more problem...

OpenStudy (anonymous):

Your personal tutor is here, at your service...

OpenStudy (lilai3):

What is the probability of spinning a number greater than 5 on a spinner numbered 1 to 8 and tossing a tail on a coin?

OpenStudy (anonymous):

the probability for an event is the amount of events you want divided by the total amount of events possible. so you can spin 1 2 3 4 5 6 7 8, of those 6 and 7 are greater. then two out of 8 are the events you want so it's 2/8 or 1/4th probability. a coin has only two outcomes, heads or tails so using the same, the probabilty for tails is 1/2

OpenStudy (anonymous):

These are independent events (the second event happens independent of the first) so we find the probability of the events and multiply. Getting a number greater than five is a probability of three out of eight (6,7, or 8). The tails on the coin is 1/2 for a fair coin. Multiply to get three of sixteen.

OpenStudy (anonymous):

woops made a mistake should obviously be 6 7 and 8 rofl

OpenStudy (anonymous):

Gotta love the importance of attention to detail....

OpenStudy (anonymous):

u mad

OpenStudy (lilai3):

thanks again so much guys!

OpenStudy (anonymous):

Not mad here. See you later.

OpenStudy (lilai3):

c u